OpenStudy (anonymous):

Find the center and radius of the circle whose equation is x^2−1x+y^2−7y−14=0.

OpenStudy (tkhunny):

You must get this idea in your head: \((a+b)^{2} = a^{2} + 2ab + b^{2}\). The really important parts are anything with a 'b' in it. This is what we leverage to solve the problem.

OpenStudy (tkhunny):

Once we identify 'x' in our problem statement with 'a' in our fundamental idea, we can focus on the values that correspond to 'b'. \(x^{2} - x\) -- We need to identify -1 as \(2b\). Given \(2b\), how do we find \(b^{2}\)?

OpenStudy (anonymous):

is it y?

OpenStudy (anonymous):

tkhunny, are you there?

OpenStudy (tkhunny):

"y"? Where did that come from? We're only talking about 'x'.

OpenStudy (tkhunny):

It is an independent question. Put everything else out of your mind. If \(2b = -1\), what is \(b^{2}\)?

OpenStudy (anonymous):

.25

OpenStudy (tkhunny):

\((-1/2)^{2} = 1/4\) - Perfect. Now to our problem statement... x^2−1x+y^2−7y−14=0 Thinking about just he x-parts, we already have the solution we need. x^2 − x + 1/4 + y^2 − 7y − 14 = 0 What do you thnk about that?

OpenStudy (anonymous):

looks similar to what i have here

OpenStudy (tkhunny):

Well that's good and bad. It is TOWARD what we want, but as stated it is horribly wrong! We can't just add 1/4 and have the same problem! We have to fix that. x^2 − x + 1/4 + y^2 − 7y − 14 = 1/4 Okay, now I feel better. We added 1/4 to each side and retained the equality. We calculated 1/4 as what we needed, and then just copied it on the other side so the world would not end. Awesome. Now, you do the same thing for the y-values. What shall we add to both sides so that the y-parts make a perfect square trinomial?

OpenStudy (anonymous):

-14

OpenStudy (tkhunny):

Bzzzzzt! (That was the bad buzzer!) You didn't ask the right question. If \(2b = -7\), what is \(b^{2}\). It is \(b^{2}\) that we need.

OpenStudy (anonymous):

12.15

OpenStudy (tkhunny):

12.25, I hope you mean. \(\left(\dfrac{7}{2}\right)^{2} = \dfrac{49}{4}\) Now we have: x^2 − x + 1/4 + y^2 − 7y + 49/4 − 14 = 1/4 + 49/4 What's next? I'm pretty excited. I think we're close!

OpenStudy (anonymous):

do we have to factor and simplify?

OpenStudy (tkhunny):

Sort of. Why did we calculate those goofy little numbers? What was our purpose in doing that? (Please refer to the fundamental idea where we started this conversation.)

OpenStudy (anonymous):

determine b? im really struggling with this sorry

OpenStudy (tkhunny):

No worries. There's a lot going on. We calculated those little numbers for the sole purpose of creating a perfect square trinomial. Given that, we can rewrite as a squared binomial. x^2 − x + 1/4 + y^2 − 7y + 49/4 − 14 = 1/4 + 49/4 Rewrite \((x - 1/2)^{2} + (y - 7/2)^{2} - 14 = 1/4 + 49/4\) Does that look better?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i figured it out thanks

OpenStudy (tkhunny):

What did you get? Hopefully, there was some simplifying going on, there.

OpenStudy (anonymous):

center of circle is 1/2, 3.5 radius is 5.14 im going to post a new question ok please look at it thanks

OpenStudy (tkhunny):

Ugly decimals. But GREAT work, otherwise.